# Reconciling Different Theories of Learning with an Agent-based Model of Procedural Learning

### v1

@@ Abstract
Agent-based modeling, ICAP, KLI, cognitive load theory, simulation
1 Introduction
@@
The ICAP framework [2] is one of the most influential recent theories of active learning as it is the most cited paper in Educational Psychologist since the date it was published. In this framework, Chi and Wylie suggest that there are four different levels of cognitive engagement (passive, active, constructive, and interactive) and that higher modes of engagement result in greater learning outcomes. While research has shown that the ICAP hypothesis holds in a variety of studies, there are several limitations to the ICAP framework.
@@
First, although it is not made explicit, ICAP primarily makes sense in the context of conceptual learning, especially when it comes to constructive and interactive learning. The authors do describe how self-explanation could be a constructive activity in the context of procedural learning; but even then, what is learned from self-explanation is “rationales and justifications,” a form of conceptual understanding [2].
@@
Second, related to the first point, there are many kinds of learning where we expect the ICAP hypothesis to not apply. This was admitted in the original ICAP paper: “for domains or topics in which students cannot bootstrap themselves into deeper understanding due to a lack of relevant schemas, or for topics for which no deeper rationales exist, the predictions made by ICAP may not hold” [2]. In particular, the authors describe how learning English articles (with somewhat arbitrary rules) could be one such domain. More generally, the Knowledge-Learning-Instruction (KLI) framework provides a formal taxonomy of knowledge components (KCs), only some of which are best learned using “understanding and sense-making processes” [9]. Therefore, it seems as though ICAP cannot account for many kinds of realistic learning scenarios. While KLI might give guidelines for what these scenarios are, these two frameworks have not been juxtaposed in an easy-to-interpret way to our knowledge.
@@
Finally, some theoretical frameworks, such as cognitive load theory (CLT) [14], actually make predictions that oppose the ICAP hypothesis. As mentioned by the authors [2]:
@@
ICAP’s predictions seem to be the opposite of predictions from the load theory. For example, cognitive load theory states that the greater the load in the to-be-processed presented materials, the more difficult a task becomes, and therefore the less resulting learning. However, ICAP makes the opposite prediction.
@@
We note that CLT has primarily been studied in the context of procedural learning tasks, which is possibly why the two theories make opposing predictions in some cases, but each theory has decent explanatory power in the areas they are primarily applied in (conceptual learning for ICAP and procedural learning for CLT).
@@
To help address all of these issues, we devised a novel computational model of learning, Procedural ABICAP. Computational models can play a vital role in achieving different purposes such as teacher training, learning by teaching, and generating or evaluating hypotheses about human learning [8]. Specifically, we developed an agent-based model of procedural learning drawing upon the ICAP framework. Our work builds upon recent work that proposed the ABICAP model, an agent-based model based on the ICAP framework [11]. However, given the focus of ICAP, the ABICAP model was also primarily designed to account for conceptual learning. Moreover, this model did not account for CLT. Therefore, we sought to see if we could develop a version of the ABICAP model that could account for (1) learning procedural knowledge, (2) the predictions made by KLI, and (3) the predictions made by CLT.
@@
2 Theoretical Frameworks
2.1 ICAP
According to ICAP, in the passive mode, students passively store the information that they are being taught about and thus gain minimal understanding. For instance, a passive activity could be listening to a lecture without doing anything else. Active learning is a higher mode of engagement in which the new information activates students’ prior knowledge and the student integrates the activated prior knowledge with new knowledge. For instance, when the learner takes notes during a lecture, they are actively learning new knowledge. The constructive mode is a higher mode of engagement in which the learner constructs new pieces of knowledge by integrating the new information with activated prior knowledge. For instance, a student who draws a concept map gains a deeper understanding which can potentially transfer to related topics they had not previously learned. Finally, interactive learning takes place in a group of two or three learners who are already learning constructively, for instance, by collaboratively drawing a concept map in a dyad. Based on a large body of empirical literature, ICAP hypothesizes that the following relationship holds in terms of learning outcomes:
@@ 2.2 KLI
The KLI framework articulated by Koedinger et al. [9] tries to determine instructional practices that best serve specific knowledge components and learning events. The framework tries to address opposing recommendations based on different frameworks of learning; for example, while some researchers advocate for “instruction that increases demands on students so as to produce desirable difficulties,” other researchers draw on cognitive load theory to advocate for decreasing the extraneous cognitive load on students [14]. KLI tries to address these oppositions by distinguishing between different knowledge components (KCs) and learning events to recommend optimal instructional practices for each setting.
@@
Different types of KCs are distinguished in KLI by several variables: the kind of input (constant or variable), the kind of response (constant or variable), whether the relationship between the input and output can be verbalized, and whether it has some kind of rationale. Moreover, there are three kinds of learning processes in KLI: memory and fluency-building, induction and refinement, and understanding and sense-making. The former two are generally useful for learning procedural KCs while the latter is useful for learning conceptual KCs. For instance, learning Chinese vocabulary involves learning a set of largely isolated facts that (typically) do not have a rationale [^1] While previous work has categorized Chinese vocabularies as KCs that are not rationale, it is worth mentioning that it is not always the case, as some types of rationales could be derived from pictographs, etc. For the purpose of this paper, we use these KCs as an example, while other types of KCs can be of similar types, such as learning the capitals of countries (i.e., France $\rightarrow$ Paris).; “while a few radicals may have iconic value, the relationship generally has no rationale, instead being a writing convention” [9]. These kinds of KCs are best supported by memory and fluency-building activities. On the other hand, KCs that have a rationale (e.g., steps in a math problem or aspects of a scientific concept) are supported by sense-making and understanding processes. We believe the latter are the kinds of KCs that ICAP typically focuses on.
2.3 Cognitive Load Theory
@@
Cognitive load theory suggests that learning is most effective when cognitive processing is kept to a manageable level, as human working memory has limited capacity [14]. It posits that instructional designs should minimize extraneous cognitive load while optimizing intrinsic and germane load to facilitate the encoding of information into long-term memory. Two main implications of CLT are the worked example effect and the expertise reversal effect. The worked example effect states that learners with low prior knowledge can benefit more from studying worked examples rather than solving problems themselves, as it reduces extraneous cognitive load, allowing them to focus on understanding the underlying principles and processes. [13]. The expertise reversal effect expands upon this by noting that “Instructional techniques that are highly effective with inexperienced learners can lose their effectiveness and even have negative consequences when used with more experienced learners” [5]. For example, in the domain of algebra equations, it has been shown that studying worked examples is more beneficial than problem-solving for novice learners (i.e., worked example effect) whereas, for learners with sufficient prior knowledge, problem-solving is associated with better outcomes compared to studying worked examples [13]. These findings are inconsistent with ICAP; based on ICAP, we would anticipate that students who solely engage in problem-solving (active or constructive) have better learning outcomes than the ones who study worked examples (passive). While, Chi and Wylie do not directly address the worked example effect, they suggest that some of the benefit of worked examples may be due to explicit self-explanation (a constructive activity) or implicit self-explanation when having students complete some of the steps in an incomplete worked example. They also note that fading worked examples (gradually requiring students to complete more steps) has been shown to be more effective as students develop more proficiency. Clearly, this last observation suggests that the benefit of constructive activities is somewhat dependent on expertise, but they do not give a theoretical explanation for why this is from within the ICAP framework.
3 Procedural ABICAP
@@
The model is largely based on the recently-proposed open-source ABICAP model [11] with a few changes made to the model in order to (1) make it reasonable for procedural learning and (2) help reconcile across ICAP, CLT, and KLI. The first change is that the ABICAP model assumed that knowledge of each piece of knowledge is binary: a learner either learns something or does not. This may be reasonable for conceptual knowledge (e.g., where each node is a small piece of declarative knowledge connected to other declarative knowledge); however, for procedural knowledge, it seems more realistic for learning to take place gradually over time as evidenced by learning curves [9]. Moreover, this seems to be a corollary of the expertise-reversal effect: if the optimal learning activity for a particular skill depends on a learner’s degree of expertise on that skill, then that suggests mastery gradually develops over several practice opportunities.
The other two ways in which we modified the model form two major assumptions underlying Procedural ABICAP. We believe these assumptions (whether right or wrong) can enhance theoretical discourse in the learning sciences in service of reconciling different theories:
@@
When a learner practices a procedural skill constructively or interactively, they do not simultaneously learn other skills, but rather they can increase in their ability to connect the practiced skill to related skills. That is, knowledge does not immediately transfer to other related skills, but rather constructive learning acts as a form of “preparation for future learning” [1] of related skills.
@@
3.1 Knowledge
Procedural ABICAP assumes that there is a set of knowledge components modeled as nodes in a knowledge graph. Each node in the knowledge graph represents a procedural knowledge component that a student has mastery over. Also, knowledge components can be connected to each other via undirected edges. For instance, the edge that connects two nodes $i$ and $j$ is named $w_{ij}$. Each edge has a default associated weight which shows the extent of the relationship between two nodes. Edge weights may be increased in the learning process for constructive and interactive learners but will remain unchanged in the learning of active and passive learners.
@@ 3.1 Knowledge
3.2 Learning
We model learning for each mode of engagement based on the proposed knowledge change process of each mode as described in the ICAP framework [2]. In each step of our model, learners practice one node named the practicing node. Here we describe the procedure of practicing implemented in Procedural ABICAP. Practicing a specific node in the model is generally described by the following function:
$\textrm{Mastery}_{t+1,i}=\\ \sigma\left(3.5\times\left(\textrm{Mastery}_{t,i}+\delta_{\text{not passive}}\sum_{j}{\textrm{Mastery}_{t,j}*w_{ij,t}}\right)-(b_{i}+cl_{\text{mode}})\right)$ (1)
@@ 3.2 Learning
$0\leq w_{ij}\leq 1,\qquad\forall t,\forall i\in[1,N_{KCs}],\quad 0\leq\textrm{Mastery}_{t,i}\leq 1$
@@
where $\textrm{Mastery}_{t,i}$ represents the mastery over node $i$ at time step $t$, $\sigma$ is the sigmoid function, $\delta_{\text{not passive}}$ is 0 if the mode is passive and 1 otherwise, $w_{ij,t}$ is the edge weight for the edge between nodes $i$ and $j$ at time $t$, $b_{i}$ represents the difficulty of learning node $i$ (which could be due to a variety of factors, including item-specific intrinsic cognitive load), and $cl_{\text{mode}}$ represents the cognitive load for the current mode of engagement. We set $b_{i}$ to be the same for all nodes in this study, as we are not attempting to model item-specific differences and we do not expect varying item difficulties to affect our overall results. We also make $cl_{\text{mode}}$ increase monotonically as the mode of engagement increases as per Assumption 1. The sigmoid function exhibits a characteristic “S” shape, where it starts with a sharp increase, transitioning smoothly to a plateau as the input value increases, reflecting its nature as a squashing function that limits its output to values between 0 and 1.
@@
Although the general practice function underlies all of the learning procedures of the modes, there are mode-specific aspects of the learning process which further differentiate the different modes of engagement, as we describe below.
3.2.1 Passive
@@
Passive learning is described as a process of storing knowledge in an isolated way [2]. In procedural learning, it translates to gaining mastery over a node regardless of other nodes. Conversely, gaining mastery in passive nodes should not affect the other nodes. As shown in equation 1, the new mastery is only a function of the previous mastery over the same node (since $\delta_{\text{n}otpassive}=0$), the cognitive load of the mode, and the difficulty of the node. Passive has the lowest cognitive load [2, 9] compared to other modes of engagement. Learning of a typical passive learner is shown in Figure 1a, where with five practices, the learner achieves mastery over a node.
@@
(a) Passive learner’s mastery over a node through steps
(b) A typical simulation with 20 nodes indicating that ICAP assumption holds
Figure 1: Typical behavior of the model
3.2.2 Active
@@
Active learning is described as “integrating” processes where prior knowledge is activated and “new information can be assimilated into the activated schema” [2]. As shown in Equation 1, for an active learner, the integrating process is modeled by taking the prior knowledge into account by having the learning rate increased as a function of the mastery over each adjacent node multiplied by the weight of the edge that connects those nodes.
3.2.3 Constructive
Constructive learning is described as learning that invokes processes such as comparing, contrasting, analogizing, generalizing, or self-explaining that not only requires integrating prior knowledge with new information but also involves “inferring“ new knowledge that goes beyond the material presented to the learner. Although the description is best interpreted for conceptual knowledge, in the domain of procedural knowledge these kinds of processes could arise when, for instance, a learner infers the relationship between different steps of solving an algebra equation. In addition to the learning described by equation 1, we model constructive learning in Procedural ABICAP by “reinforcing” edges between the practicing node and neighboring nodes, as implied by Assumption 2. Concretely, at each step of practicing, a constructive learner increases the edge weights of at most two arbitrary edges between the practicing node and neighboring nodes by a constant amount. While the exact amount the edge weights increase should be problem-specific, we used the following rules for simplicity:
for $\text{for }k,j\text{ in neighbors}(i),\quad w_{ij,t+1}=w_{ij,t}+0.15,w_{ik,t+1}=w_{ik,t}+0.1$ (2)
@@ 3.2.3 Constructive
Note that if there are no edges between the practicing node and the other nodes, then the constructive learner cannot reinforce any edges.
3.2.4 Interactive
Interactive learners learn by “co-inferring” processes that involve both partners participating in constructive learning (as described above) and taking turns sharing the knowledge that they learned and constructed with each other. As Chi and Wylie describe it “This mutuality further benefits from opportunities [and] processes to incorporate feedback, to entertain new ideas, alternative perspectives, new directions, etc.” [2] We model practicing in the interactive mode as follows: at the beginning of the simulations, interactive learners are randomly paired in dyads of two learners, and then at each step, the peers engage in an interactive practice procedure where they share their edge weights with each other. In this procedure, at each step, an edge connected to the practicing node for each peer is randomly chosen. If the weights are equal, then the weights are increased by a specific amount similar to Equation 2. If not, then the maximum weight between the peers’ corresponding edges is assigned to both. This procedure takes place two times at each step of interactive practice.
4 Simulations and Results
In this section, we show how Procedural ABICAP can be used to account for different theoretical phenomena and simulate prior empirical studies. Although we rely on simulations of the models for most of these results, we note that many of the results (including the first one which does not rely on any simulations) are obtained by virtue of the way we designed the model.
In the following simulations, unless otherwise specified, we used the following parameters. The knowledge graphs are undirected Watts-Strogatz graphs [16] with 20 nodes ($N_{KCs}=20$), where each node has three edges on average. For each mode of engagement or experimental condition, we simulate 50 learners. Each learner practices the same node $i$ until they have $\textrm{Mastery}_{i}>0.8$, at which point they move on to the next node. In our results, we report the average “knowledge” of agents in each condition, which corresponds to the average number of KCs $i$ where $\textrm{Mastery}_{i}>0.8$ for each group. Each simulation is run for 40 steps. The setting of variables is as follows: $b_{i}=0.6,cl_{passive}=0.5,cl_{active}=0.6,cl_{constructive}=0.7,cl_{interactive}=0.8$.
Before delving into more specific simulations, we need to show the general compatibility of our model with ICAP. As illustrated in Figure 1b, after enough steps, the ICAP hypothesis does indeed hold (i.e., $I>C>A>P$).
4.1 Simple Procedural KCs
@@
Chi and Wylie admit that “one type of topic for which increased cognitive engagement may not be helpful is simple procedural domains for which the rules are arbitrary and cannot be logically deduced.” In KLI terms, this would be any KC that is non-verbal and where there is no associated rationale, such as Chinese vocabulary learning and learning when to use which English articles (an example mentioned in both the ICAP [2] and KLI [9] papers). According to KLI, this includes associations, categories, and productions (or arbitrary rules). Indeed, Wylie et al. demonstrated that when learning English articles, self-explanations do not appear to be helpful. Chi and Wylie suggest this may be because there is no rationale; Wylie et al. also explained this result “could be that the act of generating and selecting explanations added extraneous cognitive load to the task.”
@@
4.2 Expertise Reversal Effect

### v2

@@ Abstract
Keywords: Agent-based modeling, ICAP, KLI, cognitive load theory, simulation
Introduction
@@
The ICAP framework (?, ?) is one of the most influential recent theories of active learning as it is the most cited paper in Educational Psychologist since the date it was published. In this framework, Chi and Wylie suggest that there are four different levels of cognitive engagement (passive, active, constructive, and interactive) and that higher modes of engagement result in greater learning outcomes. While research has shown that the ICAP hypothesis holds in a variety of studies, there are several limitations to the ICAP framework.
@@
First, although it is not made explicit, ICAP primarily makes sense in the context of conceptual learning, especially when it comes to constructive and interactive learning. The authors do describe how self-explanation could be a constructive activity in the context of procedural learning; but even then, what is learned from self-explanation is “rationales and justifications,” a form of conceptual understanding (?, ?).
@@
Second, related to the first point, there are many kinds of learning where we expect the ICAP hypothesis to not apply. This was admitted in the original ICAP paper: “for domains or topics in which students cannot bootstrap themselves into deeper understanding due to a lack of relevant schemas, or for topics for which no deeper rationales exist, the predictions made by ICAP may not hold” (?, ?). In particular, the authors describe how learning English articles (with somewhat arbitrary rules) could be one such domain. More generally, the Knowledge-Learning-Instruction (KLI) framework provides a formal taxonomy of knowledge components (KCs), only some of which are best learned using “understanding and sense-making processes” (?, ?). Therefore, it seems as though ICAP cannot account for many kinds of realistic learning scenarios. While KLI might give guidelines for what these scenarios are, these two frameworks have not been juxtaposed in an easy-to-interpret way to our knowledge.
@@
Finally, some theoretical frameworks, such as cognitive load theory (CLT) (?, ?), actually make predictions that oppose the ICAP hypothesis. As mentioned by the authors (?, ?):
@@
ICAP’s predictions seem to be the opposite of predictions from the [cognitive] load theory. For example, cognitive load theory states that the greater the load in the to-be-processed presented materials, the more difficult a task becomes, and therefore the less resulting learning. However, ICAP makes the opposite prediction.
@@
CLT has primarily been studied in the context of procedural learning tasks, which is possibly why the two theories make opposing predictions in some cases, but each theory has decent explanatory power in the areas they are primarily applied in (conceptual learning for ICAP and procedural learning for CLT).
@@
To help address all of these issues, we devised a novel computational model of learning: Procedural ABICAP. Computational models can play a vital role in achieving different purposes such as teacher training, learning by teaching, and generating or evaluating hypotheses about human learning (?, ?). Specifically, we developed an agent-based model of procedural learning drawing upon the ICAP framework. Our work builds upon the ABICAP model, an agent-based model rooted in ICAP (?, ?). However, given the focus of ICAP, the ABICAP model was also primarily designed to account for conceptual learning. Moreover, this model did not account for CLT. Therefore, we sought to see if we could develop a version of ABICAP that could account for (1) learning procedural knowledge, (2) the predictions made by KLI, and (3) the predictions made by CLT.
@@
Related Work
ICAP
According to ICAP, in the passive mode, students passively store the information that they are being taught about and thus gain minimal understanding. For instance, a passive activity could be listening to a lecture without doing anything else. Active learning is a higher mode of engagement in which the new information activates students’ prior knowledge and the student integrates the activated prior knowledge with new knowledge. For instance, when the learner takes notes during a lecture, they are learning actively. The constructive mode is a higher mode of engagement in which the learner constructs new knowledge that goes beyond the information given by integrating the new information with activated prior knowledge. For instance, a student who draws a concept map gains a deeper understanding which can potentially transfer to related topics they had not previously learned. Finally, interactive learning takes place when a group of two or more learners who are already learning constructively build on each others’ knowledge through interaction. For instance, this could be when collaboratively drawing a concept map in a dyad. Based on a large body of empirical literature, ICAP hypothesizes that the following relationship holds in terms of learning outcomes:
@@
The KLI framework articulated by Koedinger et al. (?, ?) tries to determine instructional practices that best serve specific knowledge components and learning events. The framework tries to address opposing recommendations based on different frameworks of learning; for example, while some researchers advocate for “instruction that increases demands on students so as to produce desirable difficulties,” other researchers draw on cognitive load theory to advocate for decreasing the extraneous cognitive load on students (?, ?). KLI tries to address these oppositions by distinguishing between different types of knowledge components (KCs) and learning events to recommend optimal instructional practices for each setting.
@@
Different types of KCs are distinguished in KLI by several variables: the kind of input (constant or variable), the kind of response (constant or variable), whether the relationship between the input and output can be verbalized, and whether it has some kind of rationale. Moreover, there are three kinds of learning processes in KLI: memory and fluency-building, induction and refinement, and understanding and sense-making. The first two are generally useful for learning procedural KCs while the third one is useful for learning conceptual KCs. For instance, learning Chinese vocabulary involves learning a set of largely isolated facts that (typically) do not have a rationale [^1] While previous work has categorized Chinese vocabularies as KCs that are not rationale, it is worth mentioning that it is not always the case, as some types of rationales could be derived from pictographs, etc. For the purpose of this paper, we use these KCs as an example, while other types of KCs can be of similar types, such as learning the capitals of countries (i.e., France $\rightarrow$ Paris).; “while a few radicals may have iconic value, the relationship generally has no rationale, instead being a writing convention” (?, ?). These kinds of KCs are best supported by memory and fluency-building activities. On the other hand, KCs that have a rationale (e.g., steps in a math problem or aspects of a scientific concept) are supported by understanding and sense-making processes. We believe the latter are the kinds of KCs that ICAP typically focuses on.
Cognitive Load Theory
@@
Cognitive load theory suggests that learning is most effective when cognitive processing is kept to a manageable level, as human working memory has limited capacity (?, ?). It posits that instructional designs should minimize extraneous cognitive load to facilitate the encoding of information into long-term memory. Two main implications of CLT are the worked example effect and the expertise reversal effect. The worked example effect states that learners with low prior knowledge can benefit more from studying worked examples rather than solving problems themselves, as it reduces extraneous cognitive load, allowing them to focus on understanding the underlying principles and processes (?, ?). The expertise reversal effect expands upon this by noting that “Instructional techniques that are highly effective with inexperienced learners can lose their effectiveness and even have negative consequences when used with more experienced learners” (?, ?). For example, in the domain of algebra equations, it has been shown that studying worked examples is more beneficial than problem-solving for novice learners (i.e., worked example effect) whereas, for learners with sufficient prior knowledge, problem-solving is associated with better outcomes compared to studying worked examples (?, ?). These findings are inconsistent with ICAP; based on ICAP, we would anticipate that students who solely engage in problem-solving (active or constructive) have better learning outcomes than the ones who study worked examples (passive). While, Chi and Wylie do not directly address the worked example effect, they suggest that some of the benefits of worked examples may be due to explicit self-explanation (a constructive activity) or implicit self-explanation when having students complete some of the steps in an incomplete worked example. They also note that fading worked examples (gradually requiring students to complete more steps) has been shown to be more effective as students develop more proficiency. Clearly, this last observation suggests that the benefit of constructive activities is somewhat dependent on expertise, but they do not give a theoretical explanation for why this is from within the ICAP framework.
Procedural ABICAP
@@
The model is largely based on the recently-proposed open-source ABICAP model (?, ?) with a few changes made to the model in order to (1) make it reasonable for procedural learning and (2) help reconcile across ICAP, CLT, and KLI. The first change is that the ABICAP model assumed that knowledge of each piece of knowledge is binary: a learner either learns something or does not. This may be reasonable for conceptual knowledge (e.g., where each node is a small piece of declarative knowledge connected to other declarative knowledge); however, for procedural knowledge, it seems more realistic for learning to take place gradually over time as evidenced by learning curves (?, ?). Moreover, this seems to be a corollary of the expertise-reversal effect: if the optimal learning activity for a particular skill depends on a learner’s degree of expertise on that skill, then that suggests mastery gradually develops over several practice opportunities.
The other two ways in which we modified the model form two major assumptions underlying Procedural ABICAP. We believe these assumptions (whether right or wrong) can enhance theoretical discourse in the cognitive sciences in service of reconciling different theories:
@@
When a learner practices a procedural skill constructively or interactively, they do not simultaneously learn other skills, but rather they can increase in their ability to connect the practiced skill to related skills. That is, knowledge does not immediately transfer to other related skills, but rather constructive learning acts as a form of “preparation for future learning” (?, ?) of related skills.
@@
We note that the particulars of our model as presented below are certainly open to debate. By presenting a concrete model, we hope to initiate productive discussions on the specific mechanisms underlying human learning. We especially note that the choice of constants is arbitrary and may need to be calibrated for particular settings. Specific constants were chosen to ensure learning operates on a reasonable time scale and to depict clear contrasts between different modes of learning on those time scales. However, for the particular analyses we focus on in this paper, changing the constants does not significantly change the qualitative nature of the results.
Knowledge
Procedural ABICAP assumes that there is a set of knowledge components modeled as nodes in a knowledge graph. Each node in the knowledge graph represents a procedural knowledge component that a student has mastery over. Also, knowledge components can be connected to each other via undirected edges. For instance, the edge that connects two nodes $i$ and $j$ is named $w_{ij}$. Each edge has a default associated weight which shows the extent of the relationship between two nodes. As per Assumption 2, edge weights may be increased in the learning process for constructive and interactive learners but will remain unchanged in the learning of active and passive learners.
@@
Learning
We model learning for each mode of engagement based on the proposed knowledge change process of each mode as described in the ICAP framework (?, ?). In each step of our model, learners practice one node named the practicing node. Practicing a specific node in the model is described by the following function:
$M_{t+1,i}=\sigma\Biggl{(}3.5\times\Bigl{(}M_{t,i}+\delta\sum_{j}M_{t,j}\,w_{ij,t}\Bigr{)}\\ -\bigl{(}b_{i}+cl_{\text{mode}}\bigr{)}\Biggr{)}$ (1)
@@
$0\leq w_{ij,t}\leq 1,\qquad\forall t,\forall i\in[1,N_{KCs}],\quad 0\leq M_{t,i}\leq 1$
@@
where $M_{t,i}$ represents the mastery over node $i$ at time step $t$, $\sigma$ is the sigmoid function, $\delta$ is 0 if the mode is passive and 1 otherwise, $w_{ij,t}$ is the edge weight for the edge between nodes $i$ and $j$ at time $t$, $b_{i}$ represents the difficulty of learning node $i$ (which could be due to a variety of factors, including item-specific intrinsic cognitive load), $N_{KC}$ is the number of nodes or knowledge components, and $cl_{\text{mode}}$ represents the cognitive load for the current mode of engagement. We set $b_{i}$ to be the same for all nodes in this study, as we are not attempting to model item-specific differences and we do not expect varying item difficulties to affect our overall results. We also make $cl_{\text{mode}}$ increase monotonically as the mode of engagement increases as per Assumption 1. The sigmoid function exhibits a characteristic “S” shape, where it starts with a sharp increase, transitioning smoothly to a plateau as the input value increases, reflecting its nature as a squashing function that limits its output to values between 0 and 1.
@@
Although Equation 1 underlies how a learner’s mastery over the practicing node increases under each of the four modes, there are mode-specific aspects of the learning process which further differentiate the different modes of engagement, as we describe below.
Passive
@@
Passive learning is described as a process of storing knowledge in an isolated way (?, ?). In procedural learning, this translates to gaining mastery over a node regardless of other nodes. Conversely, gaining mastery in passive nodes should not affect the other nodes. As shown in equation 1, the new mastery is only a function of the previous mastery over the same node (since $\delta_{\text{passive}}=0$), the cognitive load of passive learning, and the difficulty of the node. Passive learning has the lowest cognitive load (?, ?, ?) compared to other modes of engagement. Learning of a typical passive learner is shown in Figure 1, where with five practice opportunities, the learner achieves mastery over a node.
@@
Figure 1: Passive learner’s mastery over a node through steps
Figure 2: Simulation with 20 nodes that demonstrates the expertise reversal effect.
Active
@@
As shown in Equation 1, the impact of prior knowledge in active learning is modeled by having the learning rate increase as a function of the mastery over each adjacent node multiplied by the weight of the edge that connects those nodes (since $\delta_{\text{active}}=1$)
Constructive
In addition to the learning described by Equation 1, we model constructive learning in Procedural ABICAP by “reinforcing” edges between the practicing node and neighboring nodes, as implied by Assumption 2. Concretely, at each step of practicing, a constructive learner increases the edge weights of at most two arbitrary edges between the practicing node and neighboring nodes by a constant amount. When practicing node $i$, for a randomly chosen neighboring node $j$,
$w_{ij,t+1}=\min(w_{ij,t}+\Delta_{j,t+1},1)$ (2)
@@
where $\Delta_{j,t+1}\sim\textsc{uniform}(0,0.15)$. The same procedure is followed for another randomly chosen node that neighbors $i$. Note that if there are no edges between the practicing node and the other nodes, then the constructive learner cannot reinforce any edges.
Interactive
We model practicing in the interactive mode as follows. At the beginning, interactive learners are randomly paired into dyads. At each step, the partners practice node $i$ and reinforce edges as in constructive learning, except that the upper bound in how much they add to each edge weight is higher and dependent on their partner’s edge weights. For partners $A$ and $B$ (where superscripts indicate edge weights for the associated partner),
$w^{A}_{ij,t+1}=\min(w^{A}_{ij,t}+\Delta^{A}_{j,t+1},1)$ (3)
$w^{B}_{ij,t+1}=\min(w^{B}_{ij,t}+\Delta^{B}_{j,t+1},1)$ (4)
where $\Delta^{A}_{j,t+1},\Delta^{B}_{j,t+1}\sim\textsc{uniform}(0,0.05+\max(|w^{A}_{ij,t}-w^{B}_{ij,t}|,0.15))$. The same procedure is followed for another randomly chosen node that neighbors $i$. In effect, if $A$ already has an edge weight that is more than 0.15 larger than $B$’s analogous edge weight, then $A$’s edge weight can increase by a larger amount (i.e., $A$ benefits from $B$’s greater knowledge about the conceptual knowledge associated with $i$).
Simulations and Results
In this section, we show how Procedural ABICAP can be used to account for different theoretical phenomena.
Simple Procedural KCs
@@
Chi and Wylie admit that “one type of topic for which increased cognitive engagement may not be helpful is simple procedural domains for which the rules are arbitrary and cannot be logically deduced.” In KLI terms, this would be any KC that is non-verbal and where there is no associated rationale, such as Chinese vocabulary learning and learning when to use which English articles (an example mentioned in both the ICAP (?, ?) and KLI (?, ?) papers). According to KLI, this includes associations, categories, and productions (or arbitrary rules). Indeed, Wylie et al. demonstrated that when learning English articles, self-explanations do not appear to be helpful. Chi and Wylie suggest this may be because there is no rationale; Wylie et al. also explained this result “could be that the act of generating and selecting explanations added extraneous cognitive load to the task.”
@@
ICAP Hypothesis and Expertise Reversal Effect
In order to assess our model’s ability to replicate the ICAP hypothesis and the expertise reversal effect, we rely on a simple simulation as follows. For the knowledge graph, we used an undirected Watts-Strogatz graph (?, ?) with 20 nodes ($N_{KCs}=20$), where each node has three edges on average. For each mode of engagement or experimental condition, we simulate 50 learners. Each learner practices the same node $i$ until they have $\text{Mastery}_{i}>0.8$, at which point they move on to the next node. In our results, we report the average number of nodes “mastered” by agents in each condition (i.e., the average number of KCs $i$ where $M_{i}>0.8$ for each group. Each simulation is run for 60 steps. The setting of variables is as follows: $b_{i}=0.6,cl_{\text{passive}}=0.5,cl_{\text{active}}=0.65,cl_{\text{constructive}}=0.75,cl_{\text{interactive}}=0.8$.